# Large Language Model Compression with Global Rank and Sparsity Optimization

### v1

@@ Abstract
Low-rank and sparse composite approximation is a natural idea to compress Large Language Models (LLMs). However, such an idea faces two primary challenges that adversely affect the performance of existing methods. The first challenge relates to the interaction and cooperation between low-rank and sparse matrices, while the second involves determining weight allocation across different layers, as redundancy varies considerably among them. To address these challenges, we propose a novel two-stage LLM compression method with the capability of global rank and sparsity optimization. It is noteworthy that the overall optimization space is vast, making comprehensive optimization computationally prohibitive. Therefore, to reduce the optimization space, our first stage utilizes robust principal component analysis to decompose the weight matrices of LLMs into low-rank and sparse components, which span the low dimensional and sparse spaces containing the resultant low-rank and sparse matrices, respectively. In the second stage, we propose a probabilistic global optimization technique to jointly identify the low-rank and sparse structures within the above two spaces. The appealing feature of our approach is its ability to automatically detect the redundancy across different layers and to manage the interaction between the sparse and low-rank components. Extensive experimental results indicate that our method significantly surpasses state-of-the-art techniques for sparsification and composite approximation.
@@ 1 Introduction
Transformer-based large language models (LLMs) Vaswani et al. [2023], Touvron et al. [2023b], OpenAI et al. [2024] have achieved remarkable progress across natural language processing (NLP), computer vision, multimodal tasks, and scientific applications. Despite these successes, their massive parameter sizes pose critical challenges: they demand huge storage and memory footprints, incur slow inference speeds, and require substantial computational resources for training. Consequently, model compression Cheng et al. [2020], Wang et al. [2024], Zhu et al. [2024] has become an essential line of research for enabling real-world LLM deployment under stringent hardware constraints.
@@
Among compression strategies, quantization Han et al. [2015], Chee et al. [2023], Kuzmin et al. [2023] typically retains overall model structure by reducing the precision of weights, thus often preserving performance. By contrast, pruning Liu et al. [2017], Frankle and Carbin [2019], Sun et al. [2023], Frantar and Alistarh [2023] removes individual weights based on certain criteria (e.g., magnitude or importance scores). Although pruning is flexible and can yield substantial parameter savings, it may degrade performance unless combined with additional fine-tuning or distillation Sanh et al. [2020], especially in large-scale LLMs that encode extensive linguistic and factual knowledge Geva et al. [2021], Dai et al. [2022].
@@
To retain more critical information under aggressive compression, researchers have explored “low-rank plus sparse" decompositions Li et al. [2023], Ren and Zhu [2023], Han et al. [2024]. In this approach, the weight matrix is decomposed into a low-rank part that captures global correlations and a sparse part that highlights outliers or domain-specific knowledge. However, existing methods often rely on manually set singular-value thresholds, which can inadvertently discard medium-sized yet important singular values. Additionally, these methods frequently require computationally expensive backpropagation for parameter updates. While there is some interaction between the optimization of the low-rank and sparse components, the two parts are still relatively independent in their update processes. Lastly, due to the significant redundancy variations from early layers to deeper ones, how to allocate rank and sparsity across layers in a globally optimal manner remains unclear.
@@
In this paper, we address these issues via a novel two-stage compression framework tailored to LLMs. First, we apply robust principal component analysis (RPCA) Candes et al. [2009] to factor each weight matrix into strictly low-rank and sparse components, thereby reducing the otherwise huge search space into a low-dimensional subspace and a sparse subspace. Second, we introduce a probabilistic global optimization scheme that jointly determines which singular values in the low-rank component and which nonzero entries in the sparse component should be retained. This is done by assigning Bernoulli probabilities and updating them via policy gradient Williams [1992] on a small calibration set, avoiding heuristic thresholds or large-scale gradient updates. Critically, our method automatically detects the differing redundancy levels across layers and manages the interaction between low-rank and sparse parts, ensuring that vital parameters are kept while truly redundant ones are pruned away. We summarize our main contributions as follows:
@@
  • We propose a two-stage LLM compression approach that first uses RPCA to produce low-rank and sparse subspaces, then employs a Bernoulli-based global optimization for rank and sparsity selection.
@@
  • Our framework eliminates the need for manual thresholds or layerwise iterative backpropagation, offering an end-to-end scheme that adapts automatically to various layers’ redundancy characteristics.
@@
2 Related Work
2.1 Unstructured Pruning
Unstructured pruning eliminates individual weights by setting them to zero, providing fine-grained control over model sparsity. SparseGPT Frantar and Alistarh [2023] leverages second-order information to perform layer-wise pruning with minimal retraining, whereas Wanda Sun et al. [2024] combines weight magnitude with activation statistics for a more straightforward pruning strategy. While these methods achieve efficient pruning, they struggle to maintain performance at high sparsity levels and often require additional retraining Sanh et al. [2020], Renda et al. [2020]. BESA Xu et al. [2024] introduces a differentiable pruning framework that dynamically allocates sparsity across layers to minimize performance degradation, producing competitive results without requiring extensive retraining. Dynamic Sparse Training (DST) Liu et al. [2020] proposes an end-to-end sparse training method where trainable pruning thresholds dynamically adjust the sparsity level during training. Unlike post-training pruning methods, DST eliminates the need for iterative fine-tuning by continuously optimizing layer-wise sparsity using backpropagation. DST is designed for training sparse networks from scratch.
2.2 Low-Rank Plus Sparse
Low-rank matrix approximation is a widely adopted approach for reducing model size. By decomposing weight matrices into reduced-rank components via SVD, low-rank approximation preserves essential information while significantly lowering memory and computational costs Denton et al. [2014], Li et al. [2023]. Recent works such as LoSparse Li et al. [2023] combine low-rank decomposition with sparse corrections to improve compression efficiency. However, these methods typically rely on heuristic thresholds and necessitate iterative fine-tuning. LPAF Ren and Zhu [2023] extends this concept by first performing structured pruning to obtain a low-rank sparse model, followed by a sparsity-aware SVD decomposition of the weight matrices. A mixed-rank fine-tuning process then reconstructs the model, leading to a compressed structure where $W\approx AB$. While effective, these methods still require multiple post-processing steps to maintain performance. Beyond post-hoc compression techniques, some works explore model sparsification during pretraining. SLTrainHan et al. [2024] formulates the sparse-plus-low-rank objective directly in the pretraining stage, jointly learning sparse and low-rank representations—a different paradigm from post-hoc compression. Meanwhile, SLIM Mozaffari and Dehnavi [2024] proposes a one-shot quantized sparse plus low-rank framework that eschews retraining by combining symmetric quantization with saliency-based decomposition. However, SLIM primarily focuses on quantization-aware sparsity. By comparison, our approach directly compresses pre-trained models using robust principal component analysis (RPCA) to globally identify both low-rank and sparse structures, thus avoiding ad-hoc thresholds and extensive re-training.
3 Preliminary
4 Preliminaries
4.1 Low-Rank Approximation
@@
Low-rank approximation Chu et al. [2003] is a fundamental technique in matrix theory, widely used to reduce the parameter count in neural networks while preserving most of the model’s performance. In large language models (LLMs), weight matrices are typically high-dimensional and dense. By approximating a weight matrix $\mathbf{W}\in\mathbb{R}^{M\times N}$ as $\mathbf{U}\mathbf{V}^{\top}$ with rank $R\ll\min(M,N)$, we can achieve substantial reductions in storage and computational cost. Concretely, one typically uses Singular Value Decomposition (SVD) to write
$\mathbf{W}=\mathbf{U}\boldsymbol{\Sigma}\mathbf{V}^{\top},$ (1)
and then retains only the largest $R$ singular values $\boldsymbol{\Sigma}_{R}$, yielding
$\mathbf{W}\,\approx\,\mathbf{U}_{R}\,\boldsymbol{\Sigma}_{R}\,\mathbf{V}_{R}^{\top}.$ (2)
$\mathbf{W}\mathbf{x}\;\approx\;\mathbf{U}_{R}\bigl{(}\boldsymbol{\Sigma}_{R}(\mathbf{V}_{R}^{\top}\mathbf{x})\bigr{)},$
leading to efficiency gains.
4.2 Low-Rank Approximation with Sparse Corrections
To mitigate the shortcomings of purely low-rank approximation, recent methods Li et al. [2023], Ren and Zhu [2023] advocate combining a low-rank matrix with a sparse correction term. One splits the model weights as:
@@ 4.2 Low-Rank Approximation with Sparse Corrections
LoSparse Li et al. [2023], for example, first applies an SVD on $\mathbf{W}$ to obtain a low-rank component (with some rank $R$), and then prunes the residual $\mathbf{W}-\mathbf{U}\mathbf{V}^{\top}$ to form a sparse matrix $\mathbf{S}$. In practice, one must still decide the singular-value cutoff (or target rank) and the sparsity ratio for the residual. Often, additional fine-tuning is performed on the low-rank part to recover lost performance, or iterative pruning is applied to the sparse part Molchanov et al. [2019], which can be computationally expensive.
@@
5 Method
We begin by formulating the global objective of compressing LLM weights under a parameter budget (§ 5.1). We then describe our proposed approach (§ 5.2), which first uses RPCA to decompose each weight matrix into low-rank and sparse components, and subsequently prunes these components in a probabilistic manner, without heuristic thresholds or large-scale fine-tuning. Additional theoretical analysis can be found in the Appendix F.
@@ 5.1 Problem Formulation
Suppose we have $L$ layers in an LLM, each containing weight matrices $\{\mathbf{W}^{(l)}\}_{l=1}^{L}$. We seek compressed matrices $\{\tilde{\mathbf{W}}^{(l)}\}$ such that the total parameter count does not exceed a budget $K$, while minimizing a loss $\ell(\tilde{\mathbf{W}})$ measured on a small calibration set $\mathcal{D}$. Formally,
@@
subject $\text{subject to}\quad\mathrm{ParamCount}\bigl{(}\{\tilde{\mathbf{W}}^{(l)}\}\bigr{)}\;\leq\;K,$
@@ 5.2 Proposed Approach: CAP
As illustrated in Figure 2, we present an overview of our proposed method. Each weight matrix $\mathbf{W}$ in the model is decomposed into two components: a low-rank component $\mathbf{L}$, which captures dominant patterns, and a sparse component $\mathbf{S}$, which represents more fine-grained information. This representation enables the compression of the model while preserving its performance. To achieve this decomposition, we leverage RPCA. After decomposition, we selectively prune $\mathbf{L}$ and $\mathbf{S}$ by modeling their parameters using Bernoulli distributions, which are optimized via policy gradient. This process achieves adaptive pruning without relying on heuristic thresholds or global fine-tuning. In the following sections, we provide detailed explanations of our algorithm.
5.2.1 Robust Subspace Decomposition
The first stage aims to construct separated subspaces for low-rank and sparse components, addressing two critical challenges: (1) avoiding heuristic rank selection through automated subspace decomposition, and (2) establishing proper interaction mechanisms between structural patterns and sparse residuals. We achieve this through RPCA, which decomposes each weight matrix $\mathbf{W}\in\mathbb{R}^{m\times n}$ as:
$\mathbf{W}\;=\;\mathbf{L}+\mathbf{S}$ (5)
This decomposition fundamentally differs from conventional low-rank approximations by simultaneously addressing three key requirements: automatic rank determination, avoiding preset rank thresholds through nuclear norm optimization; outlier preservation, maintaining performance-critical sparse elements via $\ell_{1}$-constrained sparsity; and subspace orthogonality, ensuring the low-rank ($\mathbf{L}$) and sparse ($\mathbf{S}$) components inhabit complementary subspaces.
The unified RPCA optimization framework achieves these objectives through:
@@ 5.2.1 Robust Subspace Decomposition
$\min_{\mathbf{L},\mathbf{S}}\underbrace{\|\mathbf{L}\|_{*}}_{\text{Low-rank constraint}}+\lambda\,\underbrace{\|\mathbf{S}\|_{1}}_{\text{Sparsity constraint}}\text{subject to}\mathbf{W}=\mathbf{L}+\mathbf{S},$ (6)
Where the nuclear norm $\|\mathbf{L}\|_{*}$ (sum of singular values) automatically determines the effective rank through singular value shrinkage, while the $\ell_{1}$ penalty on $\mathbf{S}$ provides adaptive outlier selection by suppressing small perturbations while preserving significant deviations. The convex optimization structure guarantees global convergence, ensuring stable decomposition across different network layers.
We implement this through an Augmented Lagrangian Method (ALM) with Alternating Direction Method of Multipliers (ADMM) Lin et al. [2010], which iteratively updates:
$\mathbf{L}_{k+1}$ $=\arg\min_{\mathbf{L}}\ \|\mathbf{L}\|_{*}$
$\quad+\frac{\mu}{2}\left\|\mathbf{W}-\mathbf{L}-\mathbf{S}_{k}+\mu^{-1}\mathbf{Y}_{k}\right\|_{F}^{2},$ (7)
$\mathbf{S}_{k+1}$ $=\arg\min_{\mathbf{S}}\ \lambda\|\mathbf{S}\|_{1}$
$\quad+\frac{\mu}{2}\left\|\mathbf{W}-\mathbf{L}_{k+1}-\mathbf{S}+\mu^{-1}\mathbf{Y}_{k}\right\|_{F}^{2},$ (8)
@@
The $\mathbf{L}$-update employs Singular Value Thresholding (SVT) Cai et al. [2008]:
$\mathbf{L}_{k+1}=\mathbf{U}\operatorname{diag}(\operatorname{shrink}_{\mu^{-1}}(\boldsymbol{\sigma}))\mathbf{V}^{\top}$ (10)
where $\mathbf{U}\boldsymbol{\sigma}\mathbf{V}^{\top}$ is the SVD of $\mathbf{W}-\mathbf{S}_{k}+\mu^{-1}\mathbf{Y}_{k}$, with singular value shrinkage $\operatorname{shrink}_{\tau}(\sigma_{i})=\max(\sigma_{i}-\tau,0)$. The $\mathbf{S}$-update applies elementwise soft-thresholding:
$[\mathbf{S}_{k+1}]_{ij}=\operatorname{shrink}_{\lambda\mu^{-1}}([\mathbf{W}-\mathbf{L}_{k+1}+\mu^{-1}\mathbf{Y}_{k}]_{ij})$ (11)
@@
5.2.2 Learnable Probabilistic Pruning
In the second stage, we jointly refine important parameters in both spaces via a probabilistic global optimization technique. It is able to automatically detect redundancies between different layers and manage interactions between sparse and low-rank components without requiring computationally expensive back-propagation.
Having decomposed $\mathbf{W}$ into $\mathbf{L}$ (with singular values $\boldsymbol{\sigma}=(\sigma_{1},\dots,\sigma_{r})$) and a sparse part $\mathbf{S}$, we next decide which rank-$1$ directions in $\mathbf{L}$ and which entries in $\mathbf{S}$ to keep.
Each retained singular value $\sigma_{i}$ requires storing its corresponding singular vectors $\mathbf{u}_{i}\in\mathbb{R}^{m}$ and $\mathbf{v}_{i}\in\mathbb{R}^{n}$, contributing $(m+n)$ parameters. Denote the nonzero entries of $\mathbf{S}$ by $\{s_{ij}\}$. We introduce Bernoulli random variables:

### v3

@@ Abstract
Low-rank and sparse composite approximation is a natural idea to compress Large Language Models (LLMs). However, such an idea faces two primary challenges that adversely affect the performance of existing methods. The first challenge relates to the interaction and cooperation between low-rank and sparse matrices, while the second involves determining weight allocation across different layers, as redundancy varies considerably among them. To address these challenges, we propose a novel two-stage LLM compression method with the capability of global resource allocation for rank and sparsity. It is noteworthy that the optimization space is vast, making comprehensive optimization computationally prohibitive. Therefore, to reduce the optimization space, our first stage utilizes robust principal component analysis to decompose the weight matrices of LLMs into low-rank and sparse components, which span the low dimensional and sparse spaces containing the resultant low-rank and sparse matrices, respectively. In the second stage, we propose a probabilistic global allocation strategy to jointly identify the low-rank and sparse structures within the above two spaces. The appealing feature of our approach is its ability to automatically detect the redundancy across different layers and to manage the interaction between the sparse and low-rank components. Extensive experimental results indicate that our method significantly surpasses state-of-the-art techniques for sparsification and composite approximation.
@@ 1 Introduction
Transformer-based large language models (LLMs) (Vaswani et al., 2023; Touvron et al., 2023b; OpenAI et al., 2024) have achieved remarkable progress across natural language processing (NLP), computer vision, and scientific applications. Despite these successes, their massive parameter sizes pose critical challenges: they demand huge storage and memory footprints, incur slow inference, and require substantial computational resources for training. Consequently, model compression (Cheng et al., 2020; Wang et al., 2024a; Zhu et al., 2024; Zhou et al., 2025b; 2026) has become an essential line of research for enabling real-world LLM deployment under stringent hardware constraints.
@@
Among compression strategies, quantization (Han et al., 2015; Chee et al., 2023; Kuzmin et al., 2023) typically retains overall model structure by reducing the precision of weights, thus often preserving performance. By contrast, pruning (Liu et al., 2017; Frankle and Carbin, 2019; Sun et al., 2024; Frantar and Alistarh, 2023) removes individual weights based on certain criteria (e.g., magnitude or importance scores). Although pruning is flexible and can yield substantial parameter savings, it may degrade performance unless combined with additional fine-tuning or distillation (Sanh et al., 2020; Zhou et al., 2024; 2025a), especially in large-scale LLMs that encode extensive linguistic and factual knowledge (Geva et al., 2021; Dai et al., 2022; Cui et al., 2025).
@@
To retain more critical information under aggressive compression, researchers have explored “low-rank plus sparse” decompositions (Li et al., 2023; Ren and Zhu, 2023; Han et al., 2024). In this approach, the weight matrix is decomposed into a low-rank part that captures global correlations and a sparse part that highlights outliers or domain-specific knowledge. However, existing methods often rely on manually set singular-value thresholds, which can inadvertently discard medium-sized yet important singular values. Additionally, these methods typically require computationally expensive backpropagation for parameter updates. While there is some interaction between the optimization of the low-rank and sparse components, the two parts are still relatively independent in their update processes. Lastly, due to the significant redundancy variations from early layers to deeper ones, how to allocate rank and sparsity across layers in a globally coordinated manner remains unclear.
@@
In this paper, we address these issues via a novel two-stage compression framework tailored to LLMs. First, we apply robust principal component analysis (RPCA) (Candès et al., 2011) to factor each weight matrix into strictly low-rank and sparse components, thereby reducing the otherwise huge search space into a low-dimensional subspace and a sparse subspace. Second, we introduce a probabilistic global resource allocation scheme that jointly determines which singular values in the low-rank component and which nonzero entries in the sparse component should be retained. This is done by assigning Bernoulli probabilities and updating them via policy gradient (Williams, 1992) on a small calibration set, avoiding heuristic thresholds or backpropagation on the original LLM parameters. Critically, our method automatically detects the differing redundancy levels across layers and manages the interaction between low-rank and sparse parts, ensuring that vital parameters are kept while redundant ones are pruned away. We summarize our main contributions as follows:
@@
  • We propose a two-stage LLM compression approach that first uses RPCA to produce low-rank and sparse subspaces, then employs a Bernoulli-based global resource allocation for rank and sparsity selection.
@@
  • Our framework eliminates the need for manual thresholds or layerwise iterative backpropagation, offering a training-free scheme that adapts automatically to various layers’ redundancy characteristics.
@@
[…]
2 Method
2.1 Theoretical Background and Motivation
@@
Low-rank approximation is a fundamental technique in matrix theory, widely used to reduce the parameter count in neural networks while preserving model performance. In LLMs, weight matrices are typically high-dimensional and dense. By approximating a weight matrix $W\in\mathbb{R}^{m\times n}$ with rank $R\ll\min(m,n)$ using a truncated SVD, one can write
$W\approx U_{R}\Sigma_{R}V_{R}^{\top},$ (1)
where $U_{R}$ and $V_{R}$ contain the top $R$ left and right singular vectors, and $\Sigma_{R}$ is the diagonal matrix of the largest $R$ singular values. This factorization reduces the parameter count from $m\times n$ to $(m+n)\times R$, and breaks a large matrix multiplication into smaller ones, leading to significant efficiency gains.
Despite these benefits, low-rank approximation alone may be insufficient for LLM compression, especially when the singular values do not decay sharply. For example, Figure 3 in Appendix C shows the singular value spectra of two representative layers (Layer 0 and Layer 31) from a Transformer model, comparing the original weight matrix and its low-rank component after RPCA processing. The dashed lines (original matrices) indicate that certain modules in the same Transformer block (e.g., an attention head vs. a feed-forward network) can exhibit similar spectral shapes; yet across different layers, the redundancy patterns vary considerably. Consequently, imposing the same target rank $R$ across all layers may prune too aggressively in some cases and insufficiently in others. This observation motivates a more flexible approach that can adapt the compression ratio per layer.
@@
Recent studies have explored combining low-rank and sparse representations to enhance compression. For instance, LoSparse (Li et al., 2023) first applies SVD on $W$ to obtain a rank-$R$ approximation, then prunes the residual $W-U_{R}\Sigma_{R}V_{R}^{\top}$ to form a sparse matrix. In practice, one must still decide the singular value cutoff (or target rank) and the sparsity ratio for the residual. Often, additional fine-tuning is performed on the low-rank part to recover lost performance, or iterative pruning is applied to the sparse part (Molchanov et al., 2019), which can be computationally expensive. A major limitation of these approaches is the reliance on manually chosen thresholds for both singular values and residual pruning. They also lack a clear mechanism to coordinate how much rank vs. sparsity each layer should receive, since different layers and modules may have different redundancy characteristics. Furthermore, when both the low-rank and sparse matrices require joint fine-tuning, the memory consumption can become large, potentially exceeding the budget.
We begin by formulating the global resource allocation objective of compressing LLM weights under a parameter budget (§ 2.2). We then describe our proposed approach (§ 2.3), which first uses RPCA to decompose each weight matrix into low-rank and sparse components, and subsequently prunes these components in a probabilistic manner, without heuristic thresholds or backpropagation through the original LLM weights. Additional theoretical analysis can be found in the Appendix E.
@@ 2.2 Problem Formulation
[…]
@@
$\text{subject to}\quad\mathrm{ParamCount}\bigl(\{\tilde{\mathbf{W}}^{(l)}\}\bigr)\;\leq\;K,$
@@ 2.3 Proposed Approach: CAP
As illustrated in Figure 1, our proposed method, CAP, follows a two-stage process. The role of Stage 1 is to decompose weights into a relatively low-rank matrix $\mathbf{L}$ and a sparse matrix $\mathbf{S}$, reducing the parameter space to manageable candidates. Stage 2 then allocates the parameter budget over these candidates to achieve the target compression ratio while preserving model performance. This principled decomposition followed by budget-aware selection avoids heuristic thresholds and expensive fine-tuning. In the following sections, we provide detailed explanations of our algorithm.
2.3.1 Stage 1: Principled Decomposition via RPCA
The first stage of our method is not designed to achieve a target compression ratio directly. Instead, its purpose is to perform a principled decomposition of each weight matrix, transforming the complex problem of pruning individual weights into a more structured one. By separating a weight matrix $\mathbf{W}\in\mathbb{R}^{m\times n}$ into a low-rank component $\mathbf{L}$ that captures global structure and a sparse component $\mathbf{S}$ that captures local, salient features, we establish a high-quality candidate pool for subsequent compression. We achieve this through RPCA, which formulates the decomposition as a convex optimization problem:
@@ 2.3.1 Stage 1: Principled Decomposition via RPCA
$\min_{\mathbf{L},\mathbf{S}}\underbrace{\|\mathbf{L}\|_{*}}_{\text{Low-rank constraint}}+\lambda\,\underbrace{\|\mathbf{S}\|_{1}}_{\text{Sparsity constraint}}\text{subject to}\mathbf{W}=\mathbf{L}+\mathbf{S}.$ (3)
The choice of this objective is theoretically motivated. The nuclear norm $\|\mathbf{L}\|_{*}$ is the tightest convex relaxation of the rank function, making it the most effective convex proxy for minimizing rank. Similarly, the $\ell_{1}$ norm $\|\mathbf{S}\|_{1}$ is the standard convex relaxation for the non-convex $\ell_{0}$ norm (sparsity), which effectively identifies significant, sparse outliers. Thus, this framework provides a principled and globally optimal separation of $\mathbf{W}$ into its underlying low-rank and sparse structures.
Crucially, the hyperparameter $\lambda$ in the RPCA objective governs the nature of this decomposition, not the final compression rate. Attempting to control sparsity by simply tuning $\lambda$ leads to unpredictable changes in the rank of $\mathbf{L}$ and often results in poor-quality decompositions, a point we analyze in detail in Appendix H. Therefore, this stage focuses solely on creating an optimal candidate pool for the subsequent budget-aware pruning. We solve Eq. equation 3 using the efficient Alternating Direction Method of Multipliers (ADMM) (Lin et al., 2010). The updates are as follows:
$\mathbf{L}_{k+1}=\arg\min_{\mathbf{L}}\ \|\mathbf{L}\|_{*}+\frac{\mu}{2}\left\|\mathbf{W}-\mathbf{L}-\mathbf{S}_{k}+\mu^{-1}\mathbf{Y}_{k}\right\|_{F}^{2},$ (4)
$\mathbf{S}_{k+1}=\arg\min_{\mathbf{S}}\ \lambda\|\mathbf{S}\|_{1}+\frac{\mu}{2}\left\|\mathbf{W}-\mathbf{L}_{k+1}-\mathbf{S}+\mu^{-1}\mathbf{Y}_{k}\right\|_{F}^{2},$ (5)
@@
The $\mathbf{L}$-update employs Singular Value Thresholding (SVT) (Cai et al., 2008):
$\mathbf{L}_{k+1}=\mathbf{U}\operatorname{diag}(\operatorname{shrink}_{\mu^{-1}}(\bm{\sigma}))\mathbf{V}^{\top}$ (7)
where $\mathbf{U}\bm{\sigma}\mathbf{V}^{\top}$ is the SVD of $\mathbf{W}-\mathbf{S}_{k}+\mu^{-1}\mathbf{Y}_{k}$, with singular value shrinkage $\operatorname{shrink}_{\tau}(\sigma_{i})=\max(\sigma_{i}-\tau,0)$. The $\mathbf{S}$-update applies elementwise soft-thresholding:
$\mathbf{S}_{k+1}]_{ij}=\operatorname{shrink}_{\lambda\mu^{-1}}([\mathbf{W}-\mathbf{L}_{k+1}+\mu^{-1}\mathbf{Y}_{k}]_{ij})$ (8)
@@
2.3.2 Stage 2: Learnable Probabilistic Pruning
While the RPCA decomposition in Stage 1 provides a high-quality separation of components, it does not enforce a specific parameter budget. The second stage directly addresses this by performing a global, budget-aware selection from the candidate pools ($\mathbf{L}$ and $\mathbf{S}$) generated previously. We decide which rank-1 components in $\mathbf{L}$ and which non-zero entries in $\mathbf{S}$ to keep, to meet a user-defined parameter budget $K$ while minimizing task performance degradation.
The total parameter budget, $K$, is a user-defined hyperparameter (e.g., 50% of the original model’s parameters). Each retained singular value $\sigma_{i}$ from $\mathbf{L}$ requires storing its corresponding singular vectors $\mathbf{u}_{i}\in\mathbb{R}^{m}$ and $\mathbf{v}_{i}\in\mathbb{R}^{n}$, contributing $(m+n)$ parameters. Each retained non-zero entry of $\mathbf{S}$ contributes one parameter. We introduce Bernoulli random variables to model the retention decision for each potential parameter: